find two numbers for which the sum is 93 and the difference is 9.

let x=___________________ y=__________________
two equations:______________________________________________________________-________________________________________________________________________

Answer by stanbon(75887)

find two numbers for which the sum is 93 and the difference is 9.
-----------------------
Equations:
x + y = 93
x - y = 9
------------
Add and solve for "x":
2x = 102
x = 51
----
Solve for "y":
x + y = 93
51 + y = 93
y = 42
---------------
Cheers,
Stan H.